area of triangle is 81 cm square the measure of height is half of the base what is the measure of triangle base
Answer by edjones(8007)

a=b/2
A=ab/2=81
(b/2 * b)/2 = 81
b^2/2=81
b^2=162
b=9sqrt(2)
.
Ed